Approximately how many nucleons are there in a 1 kg object
1
Expert's answer
2020-11-30T14:51:41-0500

Mass of object μ = 1 kg

Mass of neutron "m_n = 1.67493 \times 10^{-27} \;kg"

Since the mass of nucleon = mass of neutron ≈ mass of proton

The number of nucleons

"n = \frac{μ}{m_n} \\

n = \frac{1}{1.67493 \times 10^{-27}} \\

n = 5.97 \times 10^{26}"